OpenStudy (anonymous):

What is the value of the expression when x = –26?

OpenStudy (caozeyuan):

what is the exp?

OpenStudy (anonymous):

x-6/4

OpenStudy (caozeyuan):

-26-6=-32,-32/4=-8

TheSmartOne (thesmartone):

@arrmatyphon you need to basically plug in x = -26 so \(\sf\Large \frac{-26-6}{4}\) can you tell me what -26 - 6 = ?? Hint: -26 - 6 = -1( 26 + 6) So, first what does 26 + 6 = ?? and then you just multiply -1 to that :)